OpenStudy (anonymous):

can someone help me with how to find a number that makes a and b or k so that f is continuous at every point. This is just a general question. No specifc problem.

OpenStudy (anonymous):

take the limit, replace \(k\) by that number

OpenStudy (anonymous):

the question is a little to general to answer precisely, and example would be much more enlightening

OpenStudy (anonymous):

damn that was wrong, i meant set \[k\times 4+2=16\] get \[k=\frac{7}{2}\]

OpenStudy (anonymous):

oh okay i see. thank you so much

OpenStudy (anonymous):

let me start again \[[f(x) = \left\{ \begin{array}{lr} x^2 & : x \leq 4\\ kx+2 & : x >4 \end{array} \right.\] for the first one you get \(4^2=16\) for the second one you get \[k\times 4+2\] so set \[4k+2=16\] solve for \(k\) and get \(k=\frac{7}{2}\)

OpenStudy (anonymous):

so you plug in the top and use that to get the bottom. but it would vary with the different problems

OpenStudy (anonymous):

post a different problem and we can solve it, the question is a bit vague

OpenStudy (anonymous):

okay if I have f(x)= x^2 if x is less than or equal to 4 and x+k if x is greater than 4 So, for the top I would again get 4^2=16 then, 4+k=16 solve for k and get k=4

OpenStudy (anonymous):

if i solve \(4+k=16\) i get \(k=12\)

OpenStudy (anonymous):

but that is the right idea, yes

OpenStudy (anonymous):

sorry yeah i divided by mistake

OpenStudy (anonymous):

oh okay thanks so much. I really appriciate it.

OpenStudy (anonymous):

yw